OpenStudy (anonymous):

Will fan & medal if i get help ASAP!! Write the solution in lowest terms. -1/3 + 3/4 =

OpenStudy (aum):

\[ \frac ab + \frac cd = \frac{ad+bc}{bd} \]

OpenStudy (anonymous):

????

OpenStudy (aum):

What is the least common denominator of 3 and 4?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@aum

OpenStudy (aum):

least common denominator of 3 and 4 is 12. make the denominator the same value of 12 by multiplying the top and bottom of the first fraction by 4 and the second fraction by 3.

OpenStudy (aum):

\[ -\frac 13 + \frac 34 = -\frac 13 * \frac 44 + \frac 34 * \frac 33 = -\frac{4}{12} + \frac{9}{12} = \frac{-4+9}{12} = \frac{5}{12} \]